# Let the Trial Begin: A Mock-Court Approach to Vulnerability Detection using LLM-Based Agents

### v1

@@ 3.5.3. Answering RQ3:
We experimented with three values of $k$ (i.e., the number of interaction rounds): 1, 2, and 3 inVulTrial. We selected these values to systematically investigate the impact of multiple interaction rounds while maintaining computational feasibility. For the base model, we used GPT-4o with instruction tuned agent, as it yielded the strongest overall performance.
@@ 3.5.4. Answering RQ4:
In this research question (RQ), we conduct an ablation study by systematically removing each agent to assess its necessity. Since the security researcher agent is fundamental for identifying potential vulnerabilities, we retain it in all scenarios. Additionally, since the moderator primarily provides summaries rather than making explicit decisions on vulnerabilities, we do not include a scenario where the moderator is responsible for delivering the final verdict.
In total, we evaluate four scenarios. First, the Full VulTrial (baseline) includes all four agents. Second, we exclude the moderator and use only the security researcher, code author, and review board. Third, we exclude the code author. Finally, we consider the effect of removing two agents at once. In this fourth scenario, we explore (a) only the security researcher and review board (i.e., no code author, no moderator), and (b) only the security researcher and code author (i.e., no moderator, no review board). For scenario 4(b), as the final decision is made by the code author agent, we classify the code as vulnerable if the code author mitigates at least one of the potential vulnerabilities.
@@
Table 1. Results of various LLMs on PrimeVul paired functions. An upward-pointing arrow ($\uparrow$) indicates that a higher score is better, while a downward-pointing arrow ($\downarrow$) indicates that a lower score is better.
Model Agents Method P-C $\uparrow$ P-V $\downarrow$ P-B $\downarrow$ P-R $\downarrow$
Ding et al.’s CoT (Ding et al., 2025b) 6.21 4.79 83.51 5.5
Single-Agent Finetuned 1.24 5.32 90.96 2.48
GPTLens (Hu et al., 2023) 4.60 89.20 0.69 5.51
GPT-3.5 Multi-Agent VulTrial 15.63 9.20 60.92 14.25
Single-Agent Ding et al.’s CoT (Ding et al., 2025b) 9.20 9.89 74.26 6.67
GPTLens (Hu et al., 2023) 10.11 55.40 28.05 6.44
GPT-4o Multi-Agent VulTrial 18.62 41.15 28.74 11.49
  • $\dagger$ Results taken from previous study by Ding et al. (Ding et al., 2025b).
@@ 3.6. Implementation Details
We built VulTrial on top of AgentVerse (Chen et al., 2024), a recently developed framework that simplifies the deployment of multiple LLM-based agents. For single-agent baselines, we used the prompt in PrimeVul (Ding et al., 2025b), while for GPTLens (Hu et al., 2023), we employed the provided replication package (Hu et al., 2022). Because the original GPTLens study focused on smart contracts, we adapted its prompt by replacing references to “smart contract” with “code” to generalize the use case. We used the best settings reported in their study, including two auditor agents, a top-$k$ value of 3, and the criteria for vulnerable with a final score of at least 5 (out of 9). All GPT-based models were accessed through the OpenAI API. For instruction tuning GPT-3.5 and GPT-4o, we utilized OpenAI’s official fine-tuning tools (OpenAI, 2024a). To make the results more deterministic, we set the temperature to 0 for both our single-agent baselines and VulTrial. For GPTLens (Hu et al., 2023), we followed the same setting used in its evaluation (0.7 for the auditor agent and 0 for the critic agent).
@@ 4. Results
4.1. RQ1: VulTrial vs. Single-Agent and Multi-Agent Baselines
In this RQ, we evaluated VulTrial against single-agent (using Ding et al. CoT prompts (Ding et al., 2025b)) and multi-agent (GPTLens (Hu et al., 2023)) baselines across several LLMs. We utilized GPT-3.5 and GPT-4o for the base model in VulTrial, single-agent, and multi-agent baselines (see Section 3). Table 1 presents the results, with P-C indicating pairs classified correctly for both functions (higher is better).
GPT-3.5.
Among all models, GPT-3.5 saw the largest boost when applying VulTrial. Its P-C score increased by 151.69%, over the single-agent baseline. Likewise, P-B dropped from 83.51 to 60.92, which highlights the number of both codes in the pair that were predicted as benign are less in VulTrial. Furthermore, VulTrial also improves upon the performance of GPTLens by 239.78% on P-C. In most cases, GPTLens predicts functions as vulnerable, as shown P-V score (i.e., pairs incorrectly labeled as vulnerable) of 89.2, which is much higher compared to VulTrial (9.2).
@@
GPTLens is configured to output the top-3 potential vulnerabilities per function, following the best setup from the original study (Hu et al., 2023). Since auditor agent runs twice, this results in up to six potential vulnerabilities per function. Indeed, most auditor outputs list six vulnerabilities, with an average of 5.6 and a median of 6. Meanwhile, VulTrial produce an average of 1.74 vulnerabilities per function, with a median of 2. In other words, VulTrial produces fewer vulnerabilities compared to GPTLens, but it has more precise vulnerability detection, shown by a higher P-C.
GPT-4o.
Using GPT-4o as the base model, VulTrial also achieves the best performance. Specifically, it outperforms the single-agent baseline by 102.39% and the multi-agent baseline by 84.17% in the P-C metric. Across all methods (VulTrial, single-agent, and multi-agent), performance under GPT-4o is higher than with GPT-3.5. This aligns with OpenAI’s claim that GPT-4o offers higher performance than GPT-3.5, as well as with previous benchmarking studies that compare these models (OpenAI, 2024b; Wang et al., 2024; Günay-polatkan and Sığırlı, 2024; Kipp, 2024).
For the single-agent baseline with GPT-4o, similar to its GPT-3.5 performance, it exhibited a high P-B score (pairs incorrectly classified as benign), suggesting a tendency to overlook vulnerabilities. However, in GPT-4o the P-B score dropped to 74.26, compared to GPT-3.5 (83.51). In contrast, VulTrial significantly reduced P-B to 28.74, a lower value than both the single-agent GPT-4o and its own performance under GPT-3.5 (60.92). This demonstrates VulTrial’s improved ability to correctly predict benign code with GPT-4o. Examining P-R (pairs where the predicted classification is reversed from the ground truth), VulTrial with GPT-4o also showed an improvement, reducing P-R by 24.02% compared to its performance with GPT-3.5.
@@
In the GPTLens experiment using GPT-4o, the first agent highlights a similar number of potential vulnerabilities as it did under GPT-3.5, with a median of 6 and a mean of 5.76. Meanwhile, VulTrial with GPT-4o reports fewer vulnerabilities (median 4, mean 4.11) compared to GPTLens. Despite highlighting fewer vulnerabilities, VulTrial achieved higher P-C and lower P-B scores than GPTLens, indicating overall better performance. This suggests that the additional agents in VulTrial, can effectively distinguish between justified and unjustified vulnerability claims, leading to more accurate final predictions.
@@
In Fig. 2, we show an example conversation from VulTrial. The code snippet under discussion is flagged for failing to validate negative indices, which allow an attacker to read arbitrary data by inputting out-of-range values in indices. The security researcher agent identifies two vulnerabilities: (1) Out-of-Bounds Read (CWE-125), which is valid, and (2) Type Handling, which is misleading since the existing code already handles unsupported types safely. After reviewing these findings, the code author agent agrees with the Out-of-Bounds Read assessment but refutes the Type Handling concern, which shows the agent can distinguish confirmed risks from exaggerated claims. The moderator agent then summarizes key points from both agents, which aim to help review board agent in making the decision. Based on these information, the review board agent concludes that the Out-of-Bounds Read is a genuine vulnerability. Meanwhile, for the Type Handling issue, the review board agrees with the code author’s argument that the existing implementation is safe. However, they worried about potential future changes that could introduce vulnerabilities as highlighted by the security researcher agent. As a result, the board classifies this issue as partially valid with a recommendation to monitor.
Answer to RQ1 Using GPT-3.5 as base LLMs, VulTrial improves P-C of single-agent and multi-agent baseline by 151.69% and 239.78%, respectively. Meanwhile, using GPT-4o as the base LLMs, VulTrial improve the baselines by 102.39% and 84.17%.
4.2. RQ2: Instruction Tuning on Role-Specific Agents
@@
From the RQ1’s results, we observed that VulTrial using GPT-4o achieve the best performance. In this RQ, we investigate whether instruction tuning on role-specific agents can further improve VulTrial’s overall performance. The results of this instruction tuning phase are shown in Table 2.
Table 2. Results of VulTrial using GPT-4o with instruction tuned agents. SR represents the security researcher agent, CA the code author agent, M the moderator agent, and RB the review board agent.
Instruction Tuning P-C $\uparrow$ P-V $\downarrow$ P-B $\downarrow$ P-R $\downarrow$
None 18.62 41.15 28.74 11.49
Security researcher (SR) 12.41 35.17 42.99 9.43
Code author (CA) 11.95 8.97 72.18 6.90
Moderator (M) 22.07 35.40 30.80 11.72
Review board (RB) 16.55 30.57 42.07 10.80
SR + CA 7.82 9.20 76.09 6.90
SR + CA + M 7.36 8.51 77.70 6.44
SR + CA + MR + RB 10.34 10.57 70.11 8.97
@@
All the single instruction tuned variants of GPT-4o in VulTrial, covering the security researcher agent, code author agent, moderator agent, and review board agent, outperformed both the single-agent and multi-agent baselines. Among these, tuning the moderator agent delivered the most substantial performance gain. Specifically, the moderator-tuned version increases the performance of the non-tuned version in VulTrial by 18.53% (P-C). Compared to the single-agent baseline, this instruction tuned version of VulTrial increases P-C by 139.89%. While for the multi-agent baseline, the improvement of the instruction tuned version of VulTrial increases P-C by 118.30%.
@@
Upon further analysis of the instruction tuned moderator agent, we observed that its summaries are more concise compared to the non-tuned version. Specifically, the tuned version of the moderator agent provides a shorter summary than the non-tuned version, reducing the summary length for security researchers by 16.03% and for code authors by 16.58%. In cases where the tuned VulTrial correctly classified code as benign but the non-tuned version did not, the security researcher’s summary was 21.31% shorter in the tuned version. Meanwhile, when the tuned VulTrial correctly classified code as vulnerable but the non-tuned version did not, the code author’s summary was 19.03% shorter. These observations suggest that the instruction tuned moderator agent can effectively distill key points from both the security researcher and code author, which help the review board agent to make more accurate final verdicts.
@@
Similarly, instruction tuning the code author agent altered its mitigation-to-refutation ratio. While the mitigation to refutation ratio of non-tuned version was 2:1, the tuned version exhibited a 1:4 ratio. In other words, the tuned code author refuted vulnerabilities far more frequently. From its tuning data, it learned to doubt potential vulnerabilities highlighted by security researcher. Although this approach can restrict over-diagnosing vulnerabilities, it also risks undervaluing genuine security concerns, as highlighted by an increase in P-B (cases where both snippets are incorrectly labeled benign). Despite outperforming single-agent baseline, this shows the delicate balance that has to be achieved by instruction tuning; While instruction tuning can sharpen an individual agent’s skills, it may introduce biases or extreme behaviors that undermine collective performance.
Combining Agent-Specific Tuning.
We also explored configurations with multiple tuned agents, such as the security researcher and code author, are used together in VulTrial. These combinations did not surpass the performance achieved by tuning only the moderator agent. For instance, when all agents were instruction tuned, the average number of identified vulnerabilities dropped from 4.11 to 2.60, and the refutation-to-mitigation ratio shifted to 1:4. This imbalance corresponded to a higher P-B rate of 70.11.
Answer to RQ2 Instruction tuning for moderator agent using a small dataset (50 paired samples) yields the most significant improvements. It enhances performance by 18.53% compared to the non-tuned version and increases P-C by 139.89% and 118.30% over the single-agent and multi-agent baselines, respectively.
4.3. RQ3: Effect of Multiple Rounds of Discussion
In the previous RQs, VulTrial used a single round of interaction among the agents. In this RQ, we investigate whether additional discussion rounds could enhance performance. The results of these multiple rounds experiments are shown in Table 3. The best result is achieved by using a single round interaction.

### v2

@@ 3.5.3. Answering RQ3:
We experimented with three values of $k$ (i.e., the number of interaction rounds): 1, 2, and 3 in VulTrial. We selected these values to systematically investigate the impact of multiple interaction rounds while maintaining computational feasibility. For the base model, we used GPT-4o with an instruction-tuned agent, as it yielded the strongest overall performance.
@@ 3.5.4. Answering RQ4:
In this RQ, we conduct an ablation study by systematically removing each agent to assess its necessity. Since the security researcher agent is fundamental for identifying potential vulnerabilities, we retain it in all scenarios. Additionally, since the moderator primarily provides summaries rather than making explicit decisions on vulnerabilities, we do not include a scenario where the moderator is responsible for delivering the final verdict.
We evaluated four scenarios: (1) the full VulTrial (baseline) includes all four agents; (2) removing the moderator; (3) removing the code author; (4) removing two agents simultaneously, using either (a) only the security researcher and review board or (b) only the security researcher and code author. For scenario 4(b), as the final decision is made by the code author agent, we classify the code as vulnerable if it mitigates at least one of the potential vulnerabilities.
@@ 3.6. Implementation Details
We built VulTrial on top of AgentVerse (chen2024agentverse), a recently developed framework that simplifies the deployment of multiple LLM-based agents. For single-agent baselines, we used the prompt in PrimeVul (ding2024vulnerability), while for GPTLens (hu2023large), we employed the provided replication package (gptlens). Because the original GPTLens study focused on smart contracts, we adapted its prompt by replacing references to “smart contract” with “code” to generalize the use case. We used the best settings reported in their study, including two auditor agents and a top-$k$ value of 3. We counted the results as vulnerable if they got a final score of at least 5 (out of 9). All GPT-based models were accessed via the OpenAI API, which was also used to generate instruction tuning data. For the fine-tuning itself, we used OpenAI’s built-in dashboard feature for custom GPT-3.5 and GPT-4o instruction tuning (openai2024gpt4o_finetuning). To make the results more deterministic, we set the temperature to 0 for both our single-agent baselines and VulTrial. For GPTLens (hu2023large), we followed the same setting used in its evaluation (0.7 for the auditor agent and 0 for the critic agent).
@@ 4. Results
4.1. RQ1: VulTrial vs. Baselines
Table 1 presents the performance of VulTrial against both single-agent (CodeBERT (feng2020codebert), CodeT5 (wang2021codet5), UniXCoder (guo2022unixcoder), LineVul (fu2022linevul), and Ding et al.’s CoT prompts (ding2024vulnerability)) and multi-agent (GPTLens (hu2023large)) baselines. In addition to the pair-wise metrics (ding2024vulnerability) (P-C, P-V, P-B, P-R, Error), we also report traditional metrics (P, R, FPR).
Table 1. Results of various vulnerability detection methods on PrimeVul paired functions. An upward-pointing arrow ($\uparrow$) indicates that a higher score is better, while a downward-pointing arrow ($\downarrow$) indicates that a lower score is better.
Model Agents Method P-C $\uparrow$ P-V $\downarrow$ P-B $\downarrow$ P-R $\downarrow$ Error $\downarrow$ P $\uparrow$ R $\uparrow$ FPR $\downarrow$
CodeBERT (feng2020codebert) Single-Agent Fine-tuned full (ding2024vulnerability) 5 37 386 7 430 0.49 0.10 0.10
Fine-tuned pair 0 435 0 0 435 0.50 1.00 1.00
CodeT5 (wang2021codet5) Single-Agent Fine-tuned full (ding2024vulnerability) 0 57 373 5 435 0.48 0.13 0.14
Fine-tuned pair 0 434 1 0 435 0.50 0.99 0.99
UniXCoder (guo2022unixcoder) Single-Agent Fine-tuned full (ding2024vulnerability) 4 26 403 2 431 0.52 0.07 0.06
Fine-tuned pair 6 412 17 0 429 0.50 0.96 0.95
LineVul (fu2022linevul) Single-Agent Fine-tuned§ (fu2022linevul) 7 89 336 3 428 0.51 0.22 0.21
Fine-tuned pair 0 434 1 0 435 0.50 0.10 1.00
Single-Agent Ding et al.’s CoT (ding2024vulnerability) 18 16 381 20 417 0.49 0.08 0.08
GPTLens (hu2023large) 20 388 3 24 415 0.50 0.94 0.95
GPT-3.5 Multi-Agent VulTrial 68 40 265 62 367 0.51 0.25 0.23
Single-Agent Ding et al.’s CoT (ding2024vulnerability) 40 43 323 29 395 0.53 0.19 0.17
GPTLens (hu2023large) 44 241 122 28 391 0.51 0.65 0.62
GPT-4o Multi-Agent VulTrial 81 179 125 50 354 0.53 0.60 0.52
  • $\dagger$ Fine-tuned on full PrimeVul train and val sets (4,862 vuln / 170,935 benign).
    $\ddagger$ Fine-tuned on paired PrimeVul train and val sets (3,789 vuln / 3,789 benign).
  • $\S$ Fine-tuned model from prior work (fu2022linevul).
Deep learning-based Method.
[…]
These results highlight the limitations of traditional DL models in capturing subtle vulnerability patterns at the function-pair level. While DL-based models are more lightweight and computationally efficient, our results demonstrate that this efficiency comes at a cost in actionable performance. For example, compared to the highest performing DL baseline, LineVul, achieves only 7 for P-C, with a modest FPR of 0.21 and a recall of 0.22. VulTrial (using GPT-3.5) achieves a P-C of 68 (nearly 10× higher) while maintaining a comparable FPR and recall of 0.23 and 0.25. With GPT-4o, VulTrial pushes the P-C even further to 81 (12× higher than LineVul), with recall reaching 0.60. Although this comes at the cost of a higher FPR (0.52), the trade-off results in vastly more correct pairwise classifications and a significantly higher yield of true vulnerabilities.
LLM-based Method (GPT-3.5).
The single-agent CoT prompt obtains modest accuracy (P-C=18, FPR=0.08) but misses most vulnerabilities (R=0.08). Meanwhile, GPTLens boosts recall to 0.94; however, it comes with a high FPR of 0.95 with a small number of correct pairs (20). VulTrial obtains 68 correct pairs, which is 3.8x the CoT baseline and 3.4x GPTLens, while keeping FPR to 0.23.
@@
GPTLens is configured to output the top-3 potential vulnerabilities per function, following the best setup from the original study (hu2023large). Since the auditor agent runs twice, this results in up to six potential vulnerabilities per function. Indeed, most auditor outputs list six vulnerabilities, with an average of 5.6 and a median of 6. Meanwhile, VulTrial produces an average of 1.74 vulnerabilities per function, with a median of 2. In other words, VulTrial produces fewer vulnerabilities compared to GPTLens, but it has more precise vulnerability detection, shown by a higher P-C.
LLM-based Method: GPT-4o.
VulTrial achieves the highest number of pair-wise correct predictions (P-C=81), outperforming both the CoT baseline (P-C=40) and GPTLens (P-C=44). This is 2.03× the performance of the single-agent baseline and 1.84× that of GPTLens. These gains highlight VulTrial’s effectiveness at correctly identifying both the vulnerable and benign functions within a pair that have subtle differences.
In terms of traditional classification metrics, VulTrial achieves a precision of 0.53, on par with the CoT baseline (0.53) and slightly above GPTLens (0.51). Its recall is 0.60, lower than GPTLens (0.65), but significantly higher than CoT (0.19). The FPR for VulTrial is 0.52, which is lower than GPTLens (0.62) but higher than CoT (0.17). This trade-off illustrates VulTrial’s balanced behavior; it identifies more true vulnerabilities than CoT while maintaining a lower FPR than GPTLens. In terms of P-B, which measures how often both functions in the pair were incorrectly classified as benign, VulTrial (P-B=125) is significantly lower than the CoT baseline (P-B=323) and comparable to GPTLens (P-B=122). Despite these improvements, VulTrial does not outperform the alternatives on every metric. GPTLens still achieves the highest recall, making it preferable in scenarios where catching as many vulnerabilities as possible is critical.
@@
In the GPTLens experiment using GPT-4o, the first agent highlights a similar number of potential vulnerabilities as it did under GPT-3.5, with a median of 6 and a mean of 5.76. Meanwhile, VulTrial with GPT-4o reports fewer vulnerabilities (median 4, mean 4.11) compared to GPTLens. Despite highlighting fewer vulnerabilities, VulTrial achieved higher P-C and lower P-B scores than GPTLens. This suggests that the additional agents in VulTrial, can effectively distinguish between justified and unjustified vulnerability claims, leading to more accurate final predictions.
@@
Our qualitative analysis further illustrates how VulTrial’s structured, role-specific dialogue improves both fidelity and precision of vulnerability detection. Each agent in VulTrial has a specific role. The security researcher agent analyzes the function to identify vulnerabilities. For example, in Fig. 2, the security researcher highlights two vulnerabilities: (1) an Out-of-Bounds Read due to unvalidated negative indices (CWE-125), and (2) a Type Handling issue. The first is a valid case, while the second reflects an overcautious concern since the existing code already handles unsupported types safely. After reviewing these findings, the code author agent agrees with the Out-of-Bounds Read assessment but refutes the Type Handling concern. The moderator agent then summarizes key points from both agents, which aim to help the review board agent in making the decision. Based on this information, the review board agent concludes that the Out-of-Bounds Read is a genuine vulnerability. Meanwhile, for the Type Handling issue, the review board agrees with the code author’s argument that the existing implementation is safe. However, it’s concerned about potential future changes that could introduce vulnerabilities, as highlighted by the security researcher agent. As a result, this issue is classified as partially valid, which is considered benign since it falls below the threshold. This example illustrates how the agents collectively separate confirmed vulnerabilities from exaggerated ones, preventing over-reporting and guiding appropriate remediation.
Another example of running VulTrial on vulnerable code can be found in Appendix H (ourreplicationpackageall). The security researcher agent flags the lack of strict checks; the code author proposes callbacks to mitigate the issue, and the review board ultimately rules it a high-severity, valid vulnerability requiring immediate fix. Here, both the security researcher and code author align, reinforcing the confidence of the final judgment. This demonstrates VulTrial’s capacity to not only support adversarial debate but also consensus when warranted.
Conversely, in the example of the benign function that can be found in Appendix I (ourreplicationpackageall), the security researcher agent initially raises a concern about certificate validation. The code author agent points out that X.509 verification is already used, which the review board agent independently confirms. The issue is ruled invalid. This case highlights VulTrial’s strength in filtering false positives as the security researcher agent’s overstatement is corrected by technical context from the code author’s agent.
In another example of benign function (Appendix J (ourreplicationpackageall)), the security researcher agent flags potential leakage via verbose logging. The code author proposes mitigations: increasing verbosity thresholds, securing log storage, and reviewing logged content. The moderator captures this exchange. After reviewing the evidence, the review board classifies the issue as partially valid with low severity (which is below the vulnerability threshold), opting to monitor rather than patch, thereby filtering out an overcautious report.
Our qualitative analysis shows that the four agents contribute different kinds of knowledge: pattern detection (security researcher), code-context refutation or mitigation (code author), factual summarisation (moderator), and balanced final judgment (review board). This minimizes errors on both ends: they surface genuine threats while filtering out already-handled or overstated concerns, reducing false negatives and false positives.
Answer to RQ1: VulTrial achieves the highest number of correct function pair predictions while maintaining a balanced precision and recall score. It more than doubles the number of fully‑correct function pairs relative to baselines, while avoiding the extreme false‑positive rates and the low recall.
4.2. RQ2: Instruction Tuning on Agents
@@
From the RQ1’s results, we observed that VulTrial using GPT-4o achieves the best performance. In this RQ, we investigate whether instruction tuning on role-specific agents can further improve VulTrial’s overall performance. The results are shown in Table 2.
@@
All the single instruction-tuned variants of GPT-4o in VulTrial, covering the security researcher agent, code author agent, moderator agent, and review board agent, outperformed both the single-agent and multi-agent baselines. Among these, tuning the moderator agent delivered the most substantial performance gain. Specifically, the moderator-tuned version increases the performance of the non-tuned version in VulTrial by 15 P-C scores. Compared to the single-agent baseline, this instruction-tuned version of VulTrial increases P-C score by 56. While for the multi-agent baseline, the improvement of the instruction-tuned version of VulTrial increases P-C score by 52.
Table 2. Results of VulTrial using GPT-4o with instruction-tuned agents: SR (security researcher), CA (code author), M (moderator), and RB (review board).
Instruction-tuned P-C $\uparrow$ P-V $\downarrow$ P-B $\downarrow$ P-R $\downarrow$
None 81 179 125 50
Security researcher (SR) 54 153 187 41
Code author (CA) 52 39 314 30
Moderator (M) 96 154 134 51
Review board (RB) 72 133 183 47
SR + CA 34 40 331 30
SR + CA + M 32 37 338 28
SR + CA + M + RB 45 46 305 39
@@
Upon further analysis of the instruction-tuned moderator agent, we observed that its summaries are more concise compared to the non-tuned version. Specifically, the tuned version of the moderator agent provides a shorter summary than the non-tuned version, reducing the summary length for security researchers by 16.03% and for code authors by 16.58%. In cases where the tuned VulTrial correctly classified code as benign but the non-tuned version did not, the security researcher’s summary was 21.31% shorter in the tuned version. Meanwhile, when the tuned VulTrial correctly classified code as vulnerable but the non-tuned version did not, the code author’s summary was 19.03% shorter. These observations suggest that the instruction-tuned moderator agent can effectively distill key points from both the security researcher and code author, which help the review board agent to make more accurate final verdicts.
@@
Similarly, instruction tuning the code author agent altered its mitigation-to-refutation ratio. While the mitigation to refutation ratio of the non-tuned version was 2:1, the tuned version exhibited a 1:4 ratio. In other words, the tuned code author refuted vulnerabilities more frequently. From its tuning data, it learned to doubt potential vulnerabilities highlighted by security researcher. Although this approach can restrict over-diagnosing vulnerabilities, it risks undervaluing genuine security concerns, as highlighted by an increase in P-B. Despite outperforming the single-agent baseline, this shows the delicate balance required for instruction tuning. While instruction tuning can sharpen an individual agent’s skills, it may introduce biases or extreme behaviors that undermine collective performance.
@@
Combining Agent-Specific Tuning.
We also explored configurations with multiple tuned agents, such as the security researcher and code author, are used together in VulTrial. These combinations did not surpass the performance achieved by tuning only the moderator agent. For instance, when all agents were instruction-tuned, the average number of identified vulnerabilities dropped from 4.11 to 2.60, and the refutation-to-mitigation ratio shifted to 1:4. This imbalance corresponded to a higher P-B (331). These lower results may stem from the instruction tuning setup, where the review board is trained on histories that include inputs from other agents. In the tuning data, the review board often learns to “follow” the code author’s position (especially strong arguments), which is aligned with the ground truth. Once the code author is instruction-tuned and becomes more assertive in refuting vulnerabilities, the review board (also fine-tuned) tends to side with the author. Future work should decouple agent training, for example, including cases where the review board correctly identifies a vulnerability despite strong refutations, to reduce over-alignment and improve decisions.
Answer to RQ2: Instruction tuning for moderator agent using 50 paired samples yields the most significant improvements. It improves P-C by 15, 56, and 52 compared to the non-tuned version, single-agent, and multi-agent baselines, respectively.
4.3. RQ3: Effect of Multi Rounds Discussion